# A lower bound for the first non-zero basic eigenvalue on a singular Riemannian foliation

### v2

@@ 1. Introduction
In the case that $\mathrm{Ric}(M)\geq(n-1)K\geq 0$ for some constant $K$, many results have been found: Lichnerowicz [Lic58] and Obata [Oba62] proved that on a closed Riemannian manifold with $K>0$, one has $\lambda_{1}\geq nK$ and the equality holds if and only if $M$ is isometric to the $n$-sphere of radius $1/\sqrt{K}$. In the case $\mathrm{Ric}(M)\geq 0$, we cannot get a positive lower bound without a restriction of the diameter $\mathrm{d}_{M}$ of $M$: for example, in the case of torus, if we increase the diameter arbitrarily and approaching infinity, then the first eigenvalue approaches 0. Moreover, the diameter restriction is automatic in the case of $K>0$ by Bonnet-Myer’s theorem. Hence, the diameter restriction is necessary. By Li and Yau [LY80] and the improvement by Li [Li79], we get a gradient estimate for the first nontrivial eigenfunction and derive that
@@
Zhong and Yang [ZY84] enhanced this result to be an optimal one for the case of $\mathrm{Ric}(M)\geq 0$:
@@
Hang and Wang [HW07] later provided and proved the rigidity statemtent for Zhong-Yang estimate: the equality holds if and only if $M$ is isometric to a circle of radius $\mathrm{d}_{M}/\pi$.
We can see there are many works about general optimal lower bound estimate for $\lambda_{1}$ for a given lower bound (n1)K(n-1)K of the Ricci curvature, for example, in [Kro92, CW97, AC13, ZW17], and see also [BQ00, AN12, He13, Oli+20]. Inspired by foundational results of Lichnerowicz and Zhong-Yang, Li had conjectured that the first positive eigenvalue should satisfy the following estimate (see [Yan99])
@@
for some constant $\alpha$ were made. A remarkable result by Shi and Zhang in [SZ07] proved that for any $s\in(0,1)$
@@
In [AC13], B. Andrews and J. Clutterbuck showed that the inequality of this form 1.1 with $\alpha=\frac{1}{2}$ is the best possible constant and this means Li’s conjecture is false. For more detail of the history, see [He13].
So how is the lower bound estimates of the first eigenvalue related to the object we will discuss in this paper, the first basic eigenvalue on a singular Riemannian foliation? We first introduce the concept of singular Riemannian foliations. A singular Riemannian foliation on a Riemannian manifold $M$, roughly speaking, is a partition of $M$ into connected injectively immersed submanifolds (we call leaves), which are locally equidistant to one another (see subsection 2.1). One class of examples of singular Riemannian foliations are orbits of isometric actions of Riemannian manifolds, that we call homogeneous. Another class of examples are connected components of the fibers of Riemannian submersions, that we call simple. For more detail about homogeneous and simple singular Riemannian foliations, see [Rad].
Given $(M,\mathcal{F})$ a singular Riemannian foliation, we define $\mathcal{F}$-basic function, is a function that is constant on each leaf of $\mathcal{F}$. When the know which $\mathcal{F}$ we talk about, we can call it briefly basic function. This allows us to define $\mathcal{F}$-basic Laplacian as an operator that sends basic smooth functions to basic smooth functions and define $\mathcal{F}$-basic eigenvalues, that we will discuss more detail in Section 2. When $(M,\mathcal{F})$ admits a basic mean curvature, it turns out that the ordinary Laplacian coincides with the basic Laplacian, and any results that work for the ordinary Laplacian can be used to derive corresponding results for the basic Laplacian. In the case of estimate a lower bound for the first non-zero basic eigenvalue $\lambda_{1}^{B}$, if we use Lichnerowiz’s estimate, we can get $\lambda_{1}^{B}\geq\lambda_{1}\geq nK$ but it does not imply any better result because $nK$ does not depend on $\mathcal{F}$ that we consider, and we do not see the difference between Lichnerowicz’s estimate for the first non-zero eigenvalue and Lichnerowic’z estimate for the first non-zero basic eigenvalue. However, for the estimates involving the diameter like in [ZY84, Kro92, CW97, AC13, ZW17, SZ07], we can generalize them by replacing the diameter $\mathrm{d}_{M}$ of the ambient manifold $M$ with the diameter of the leaf space $\mathrm{d}_{M/\mathcal{F}}$ and it is reasonable to expect similar estimates with leaf space diameters. Hence, in this paper, we assume that the Ricci curvature of the ambient manifold is bounded below by a non-negative value.
@@ 2.1.1. Basic definitions and facts
Singular Riemannian foliations, first introduced in [Mol88] by Molino in his study of Riemannian foliations and constitute nowadays an active field of research.
@@
The first condition in Definition 2.1 is equivalent to the leaves being locally equidistant, see [Rad], Proposition 2.3. If every leaf of $\mathcal{F}$ is closed, we say $\mathcal{F}$ is closed. In this case, the condition of local equidistance of leaves becomes the global equidistance of leaves.
@@
Then the natural projection map $\sigma:M\to M/\mathcal{F}$ where $M/\mathcal{F}$ is equipped the metric $\rho$ is a map so-called manifold submetry. For more detail, see [MR20]. We can define the diameter of the leaf space $M/\mathcal{F}$ by
@@ 2.1.2. Basic mean curvature
Let $(M,\mathcal{F})$ be a singular Riemannian foliation. Then there exists an open, dense, saturated, full-measure subset $N\subset M$, which is called the principal part, such that $N/\mathcal{F}\subset M/\mathcal{F}$ is a smooth Riemannian manifold, and the restriction of $\sigma$ on $N$, $\sigma|_{N}:N\to N/\mathcal{F}$ is a Riemannian submersion (see [LR18], page 3).
For every $p\in N$, denote $H(p)\in T_{p}M$ for the mean curvature vector of the leaf $L_{p}\subset M$. These mean curvature vectors form a smooth vector field $H$ on $N$. We say that $\mathcal{F}$ admits a basic mean curvature when $H$ is $\sigma$-related to a smooth vector field $\overline{H}$ on $N/\mathcal{F}$. Admitting a basic mean curvature is automatic when considering a singular Riemannian foliation on a round sphere (see [AR15], Proposition 3.1), or on a homogeneous foliation.
@@ 2.2.1. Averaging operator
When $\mathcal{F}$ admits a basic mean curvature, the operator $\mathrm{Av}$ takes smooth functions to smooth functions, and it commutes with the Laplacian on the principal part NN (see [LR18], Theorem 3.3 and Lemma 3.2). In particular, see [PR96] for more detail for the case of regular Riemannian foliations.
@@ 2.2.2. Basic Laplacian
In [PR96], we can see the definition of the basic Laplacian for Riemannian foliations.
@@ 3. Examples
In this section, we provide some examples to verify the result of Theorem 1.1 and Theorem 1.3 that will be proved in Sections 4 and 5. In particular, we consider singular Riemannian foliations on spheres, where we know that they have basic mean curvature, see [LR18].
@@
Isoparametric foliations on spheres are well-known examples of singular Riemannian foliations, where the reader can find a long history with contributions by many mathematicians, see [Dom] and [Chi21] for more detail.
Such a singular Riemannian foliation $(S^{n},\mathcal{F})$ can be determined by an isoparametric submanifold $N\subset S^{n}$ and taking parallel submanifolds of that $N$ to obtain an isoparametric foliation on that sphere $S^{n}$. Let $g$ be the number of principal curvatures of $N\subset S^{n}$. Then, in the case when $N$ is of codimension one, Münzner proved that $g\in\{1,2,3,4,6\}$ (see Satz A, [Mün81]).
Also in the proof of Theorem 2 of Münzner’s paper (see Satz 2, [Mün81]), the (homogeneous) polynomial $G=g(g+n-1)F-cr^{g}$ of degree $g$ is harmonic, where $c$ is chosen ($c=0$ if $g$ is odd), $r^{2}=x_{0}^{2}+\cdots+x_{n}^{2}$ and $F:\mathbb{R}^{n+1}\to\mathbb{R}$ is the Cartan-Münzner (homogeneous) polynomial of degree $g$ that satisfies $\Delta F=cr^{g-2}$ and $\|\nabla F\|^{2}=g^{2}r^{2g-2}$ ($\nabla$ and $\Delta$ are gradient and Laplacian on $\mathbb{R}^{n+1}$). Since the restriction of $G$ on $S^{n}$ is basic, $G|_{S^{n}}$ is a basic eigenfunction of $(S^{n},\mathcal{F})$ and hence, g(g+n1)g(g+n-1) is a basic eigenvalue of $(S^{n},\mathcal{F})$. Furthermore, the algebra of basic (homogeneous) polynomials is finitely generated by two generators $\rho_{1}=r^{2}$ and $\rho_{2}=F$. Hence the lowest degree of a basic homogeneous polynomial which restricts to a non-constant function on the sphere is $g$. This implies that $\lambda_{1}^{B}=g(g+n-1)$ is the first non-zero basic eigenvalue of $(S^{n},\mathcal{F})$.
The diameter of the leaf space $S^{n}/\mathcal{F}$ is exactly the distance between two singular leaves, which are level sets $V^{-1}(-1)$ and $V^{-1}(1)$ of the map $V=F|_{S^{n}}$, and hence we get $\mathrm{d}_{S^{n}/\mathcal{F}}=\pi/g$ (see Satz 4, [Mün81]).
@@ 4. The case of Non-negative Ricci curvature
We recall again the gradient estimate by Zhong and Yang in Proposition 4.1, see [ZY84] (also see the Theorem 2.3.2 in [DW]). Once notices that their argument works with not the first eigenvalue, but also with an arbitrary eigenvalue $\lambda>0$ in the spectrum.
@@
Remarks.  Notice that $\psi$ is an odd function which is continuous on $\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ and is the solution of the ODE
@@
See [SY97], section 44 of chapter 33, or also see [HW07].
We are now ready to prove Theorem 1.1. In the following proof, it is clear that when we have a gradient estimate, we can integrate over a geodesic joining two points of level sets of $u$ on both sides of the gradient estimate to derive an estimate for the first non-zero eigenvalue as many works in [LY80, ZY84] have done. The new thing in this proof is that we need to choose a more specific geodesic that works with the distances of leaves.
@@
First, we will recall a remarkable works of G.Xu and X.Xue that will be used in our main theorem of this section. More precisely, we will use Proposition 3.2 and a part that is modified from the statement of Theorem 4.1 (1), see [XX25]. Here we try to provide the outline of each proof, and for more detail, please visit [XX25], pages 8 - 13.
Proposition 4.2 (Proposition 3.2 (N) in [XX25]).
@@
In the next result of G.Xu and X.Xue that we want to use, the statement of Theorem 4.1 (1) in [XX25] does not seem to be correct as stated. More precisely, they claimed that with the same given conditions as in Proposition 4.2 and in addition, $\partial M=\emptyset$, then $M^{n}$ is isometric to $S^{1}\left(\frac{\mathcal{N}(u)}{2\sqrt{\lambda}}\right)\times N^{n-1}$, where $\mathcal{N}(u)$ is the number of nodal domains of $u$ and $S^{1}\left(\frac{\mathcal{N}(u)}{2\sqrt{\lambda}}\right)$ is a round circle with radius $\frac{\mathcal{N}(u)}{2\sqrt{\lambda}}$ and $\mathcal{N}(u)=2k$ for some $k\in\mathbb{Z}_{+}$. We provide a class of counterexamples for this statement: consider $M=\mathbb{R}\times_{\mathbb{Z}}N=\mathbb{R}\times N/\mathbb{Z}$ where $N^{n-1}$ is an (n1)(n-1)-dimensional compact Riemannian manifold of nonnegative Ricci curvature and the $\mathbb{Z}$-action is defined by $k\cdot(t,x)=\left(t+\frac{2k\pi}{\sqrt{\lambda}},\varphi^{k}(x)\right)$, where $k\in\mathbb{Z},(t,x)\in\mathbb{R}\times N$ and $\varphi:N\to N$ is an isometry on $N$. As mentioned earlier in Theorem 1.2, $M=\left[0,\frac{2\pi}{\sqrt{\lambda}}\right]\times N/\sim$ is the mapping torus of $\varphi$, where $(0,x)\sim\left(\frac{2\pi}{\sqrt{\lambda}},\varphi(x)\right)$. The function $u([(t,x)])=-\cos(\sqrt{\lambda}t)$ is the eigenfunction of $\lambda$ that satisfies the given conditions with the number of nodal domains $\mathcal{N}(u)=2$. In the case of a unit circle $N=S^{1}$, $\lambda=1$ and we obtain torus, Klein bottle (not a product) with the corresponding isometries $\varphi$ are the identity, the reflection respectively. We can even have many good examples of orientable manifolds that are not products by constructing mapping tori of some isometries, for example, we choose a torus $N=\mathbb{T}^{2}=S^{1}\times S^{1}=\mathbb{R}^{2}/\mathbb{Z}^{2}$ and construct a mapping torus of the isometry $\varphi=\left[\begin{matrix}-1&0\\ 0&-1\end{matrix}\right]$. This is one of the six compact, connected orientable flat Riemannian Seifert manifolds. Recall that a Seifert manifold is a closed 3-dimensional manifold together with a decomposition into a disjoint union of circles (called fibers) such that each fiber has a tubular neighborhood that forms a standard fibered torus. If the mapping torus $M_{\varphi}$ of $\varphi=\left[\begin{matrix}-1&0\\ 0&-1\end{matrix}\right]$ is a product of $S^{1}(r)\times P$ where $S^{1}(r)$ is a circle of some radius $r$ and $P$ is a compact $2$-dimensional manifold, then we imply that $P$ is flat. To make the product is orientable and connectedness, the only possibility of $P$ is to be a torus, and then, $S^{1}(r)\times P$ is also a compact, connected orientable flat Riemannian Seifert manifold which can be distinguished with $M_{\varphi}$ as classified. For more detail of the classification, see Theorem 2, section 8.2 in [Orl72].
@@
Therefore, in the following theorem, we recall the given conditions from Theorem 4.1(1) of [XX25] with a modified conclusion and a modified argument for Step 3 of the proof of Theorem 4.1(1) in [XX25].
Theorem 4.3 (A modification of Theorem 4.1 (1) in [XX25]).
@@
Step 1 and Step 2 (see proof of Theorem 4.1 in [XX25]): Let $N^{n-1}=u^{-1}(-1)\cap\overline{M_{p}}$, which is an (n1)(n-1)-dim hypersurface in $M^{n}$ by Proposition 4.2 (if $u^{-1}(-1)\cap\overline{M_{p}}=\emptyset$, we define $N^{n-1}=u^{-1}(1)\cap\overline{M_{p}}$, similar argument applies); and $U_{\varepsilon}(N^{n-1})=\{x\in M^{n}:\mathrm{d}(x,N^{n-1})<\varepsilon\}$. Then we get a local chart $\{x_{1},x^{\prime}\}$ for $U_{r_{0}}(N^{n-1})$, where $x^{\prime}=(x_{2},\cdots,x_{n})$ is a coordinate chart for $N^{n-1}$ and
@@
Step 3: As claimed in [XX25], for any $q\in N^{n-1}$, we can find $z\in U_{r_{0}}^{-}(N^{n-1})$ such that $F(z)=F(q)=0$ and $F(z)=\max_{x\in\overline{U_{r_{0}}^{-}(N^{n-1})}}F(x)=0$. Note $|\nabla u(z)|>0,|u(z)|<1$ and $|\nabla\sin^{-1}u(z)|^{2}=\lambda$; from Proposition 4.2, we get the local splitting domain on $M_{z}$, a component of $\{x\in M:|\nabla u(x)|>0\}$ containing $z$.
@@
Remarks.  In the case when $\mathcal{F}$ is given as the (connected) fibers of a Riemannian submersion $\pi:M\to M/\mathcal{F}$ with totally geodesic (or, more generally, minimal) fibers, if we apply Hang-Wang rigidity for the first non-zero eigenvalue of $M/\mathcal{F}$ which is also the first non-zero basic eigenvalue of $M$, then we can only conclude that $M/\mathcal{F}$ is isometric to $S^{1}\left(\frac{1}{\sqrt{\lambda_{1}^{B}}}\right)$ without knowing the structure of $\mathcal{F}$. By proving Theorem 1.2, not just know $M/\mathcal{F}$, but we also understand the shape of $M$ and the structure of $\mathcal{F}$.

### v3

@@ 1. Introduction
In the case that $\mathrm{Ric}(M)\geq(n-1)K\geq 0$ for some constant $K$, many results have been found: Lichnerowicz [13] and Obata [19] proved that on a closed Riemannian manifold with $K>0$, one has $\lambda_{1}\geq nK$ and the equality holds if and only if $M$ is isometric to the $n$-sphere of radius $1/\sqrt{K}$. In the case $\mathrm{Ric}(M)\geq 0$, we cannot get a positive lower bound without a restriction of the diameter $\mathrm{d}_{M}$ of $M$: for example, in the case of torus, if we increase the diameter arbitrarily and approaching infinity, then the first eigenvalue approaches 0. Moreover, the diameter restriction is automatic in the case of $K>0$ by Bonnet-Myer’s theorem. Hence, the diameter restriction is necessary. By Li and Yau [11] and the improvement by Li [12], we get a gradient estimate for the first nontrivial eigenfunction and derive that
@@
Zhong and Yang [29] enhanced this result to be an optimal one for the case of $\mathrm{Ric}(M)\geq 0$:
@@
Hang and Wang [8] later provided and proved the rigidity statemtent for Zhong-Yang estimate: the equality holds if and only if $M$ is isometric to a circle of radius $\mathrm{d}_{M}/\pi$.
We can see there are many works about general optimal lower bound estimate for $\lambda_{1}$ for a given lower bound (n1)K(n-1)K of the Ricci curvature, for example, in [10, 4, 1, 28], and see also [3, 2, 9, 20]. Inspired by foundational results of Lichnerowicz and Zhong-Yang, Li had conjectured that the first positive eigenvalue should satisfy the following estimate (see [27])
@@
for some constant $\alpha$ were made. A remarkable result by Shi and Zhang in [25] proved that for any $s\in(0,1)$
@@
In [1], B. Andrews and J. Clutterbuck showed that the inequality of this form 1.1 with $\alpha=\frac{1}{2}$ is the best possible constant and this means Li’s conjecture is false. For more detail of the history, see [9].
So how is the lower bound estimates of the first eigenvalue related to the object we will discuss in this paper, the first basic eigenvalue on a singular Riemannian foliation? We first introduce the concept of singular Riemannian foliations. A singular Riemannian foliation on a Riemannian manifold $M$, roughly speaking, is a partition of $M$ into connected injectively immersed submanifolds (we call leaves), which are locally equidistant to one another (see subsection 2.1). One class of examples of singular Riemannian foliations are orbits of isometric actions of Riemannian manifolds, that we call homogeneous. Another class of examples are connected components of the fibers of Riemannian submersions, that we call simple. For more detail about homogeneous and simple singular Riemannian foliations, see [23].
Given $(M,\mathcal{F})$ a singular Riemannian foliation, we define $\mathcal{F}$-basic function, is a function that is constant on each leaf of $\mathcal{F}$. When the know which $\mathcal{F}$ we talk about, we can call it briefly basic function. This allows us to define $\mathcal{F}$-basic Laplacian as an operator that sends basic smooth functions to basic smooth functions and define $\mathcal{F}$-basic eigenvalues, that we will discuss more detail in Section 2. When $(M,\mathcal{F})$ admits a basic mean curvature, it turns out that the ordinary Laplacian coincides with the basic Laplacian, and any results that work for the ordinary Laplacian can be used to derive corresponding results for the basic Laplacian. In the case of estimate a lower bound for the first non-zero basic eigenvalue $\lambda_{1}^{B}$, if we use Lichnerowiz’s estimate, we can get $\lambda_{1}^{B}\geq\lambda_{1}\geq nK$ but it does not imply any better result because $nK$ does not depend on $\mathcal{F}$ that we consider, and we do not see the difference between Lichnerowicz’s estimate for the first non-zero eigenvalue and Lichnerowic’z estimate for the first non-zero basic eigenvalue. However, for the estimates involving the diameter like in [29, 10, 4, 1, 28, 25], we can generalize them by replacing the diameter $\mathrm{d}_{M}$ of the ambient manifold $M$ with the diameter of the leaf space $\mathrm{d}_{M/\mathcal{F}}$ and it is reasonable to expect similar estimates with leaf space diameters. Hence, in this paper, we assume that the Ricci curvature of the ambient manifold is bounded below by a non-negative value.
@@
Remark 1.2.
Particularly, in the case of small diameter of $M$, we can choose a specific value of $s$ to get the optimal estimate:
$\lambda_{1}^{B}\geq\frac{\pi^{2}}{\mathrm{d}_{M/\mathcal{F}}^{2}}+\frac{1}{2}(n-1)K+\frac{(n-1)^{2}K^{2}\mathrm{d}_{M/\mathcal{F}}^{2}}{16\pi^{2}}.$
We will discuss more after proving Theorem 1.3 in Section 5.
@@ 2.1.1. Basic definitions and facts
Singular Riemannian foliations, first introduced in [17] by Molino in his study of Riemannian foliations and constitute nowadays an active field of research.
@@
The first condition in Definition 2.1 is equivalent to the leaves being locally equidistant, see [23], Proposition 2.3. If every leaf of $\mathcal{F}$ is closed, we say $\mathcal{F}$ is closed. In this case, the condition of local equidistance of leaves becomes the global equidistance of leaves.
@@
Then the natural projection map $\sigma:M\to M/\mathcal{F}$ where $M/\mathcal{F}$ is equipped the metric $\rho$ is a map so-called manifold submetry. For more detail, see [15]. We can define the diameter of the leaf space $M/\mathcal{F}$ by
@@ 2.1.2. Basic mean curvature
Let $(M,\mathcal{F})$ be a singular Riemannian foliation. Then there exists an open, dense, saturated, full-measure subset $N\subset M$, which is called the principal part, such that $N/\mathcal{F}\subset M/\mathcal{F}$ is a smooth Riemannian manifold, and the restriction of $\sigma$ on $N$, $\sigma|_{N}:N\to N/\mathcal{F}$ is a Riemannian submersion (see [14], page 3).
For every $p\in N$, denote $H(p)\in T_{p}M$ for the mean curvature vector of the leaf $L_{p}\subset M$. These mean curvature vectors form a smooth vector field $H$ on $N$. We say that $\mathcal{F}$ admits a basic mean curvature when $H$ is $\sigma$-related to a smooth vector field $\overline{H}$ on $N/\mathcal{F}$. Admitting a basic mean curvature is automatic when considering a singular Riemannian foliation on a round sphere (see [14], Proposition 3.1), or on a homogeneous foliation.
@@ 2.2.1. Averaging operator
When $\mathcal{F}$ admits a basic mean curvature, the operator $\mathrm{Av}$ takes smooth functions to smooth functions, and it commutes with the Laplacian on the principal part NN (see [14], Theorem 3.3 and Lemma 3.2). In particular, see [22] for more detail for the case of regular Riemannian foliations.
@@ 2.2.2. Basic Laplacian
In [22], we can see the definition of the basic Laplacian for Riemannian foliations.
@@ 3. Examples
In this section, we provide some examples to verify the result of Theorem 1.1 and Theorem 1.3 that will be proved in Sections 4 and 5. In particular, we consider singular Riemannian foliations on spheres, where we know that they have basic mean curvature, see [14].
@@
Isoparametric foliations on spheres are well-known examples of singular Riemannian foliations, where the reader can find a long history with contributions by many mathematicians, see [7] and [5] for more detail.
Such a singular Riemannian foliation $(S^{n},\mathcal{F})$ can be determined by an isoparametric submanifold $N\subset S^{n}$ and taking parallel submanifolds of that $N$ to obtain an isoparametric foliation on that sphere $S^{n}$. Let $g$ be the number of principal curvatures of $N\subset S^{n}$. Then, in the case when $N$ is of codimension one, Münzner proved that $g\in\{1,2,3,4,6\}$ (see Satz A, [18]).
Also in the proof of Theorem 2 of Münzner’s paper (see Satz 2, [18]), the (homogeneous) polynomial $G=g(g+n-1)F-cr^{g}$ of degree $g$ is harmonic, where $c$ is chosen ($c=0$ if $g$ is odd), $r^{2}=x_{0}^{2}+\cdots+x_{n}^{2}$ and $F:\mathbb{R}^{n+1}\to\mathbb{R}$ is the Cartan-Münzner (homogeneous) polynomial of degree $g$ that satisfies $\Delta F=cr^{g-2}$ and $\|\nabla F\|^{2}=g^{2}r^{2g-2}$ ($\nabla$ and $\Delta$ are gradient and Laplacian on $\mathbb{R}^{n+1}$). Since the restriction of $G$ on $S^{n}$ is basic, $G|_{S^{n}}$ is a basic eigenfunction of $(S^{n},\mathcal{F})$ and hence, g(g+n1)g(g+n-1) is a basic eigenvalue of $(S^{n},\mathcal{F})$. Furthermore, the algebra of basic (homogeneous) polynomials is finitely generated by two generators $\rho_{1}=r^{2}$ and $\rho_{2}=F$. Hence the lowest degree of a basic homogeneous polynomial which restricts to a non-constant function on the sphere is $g$. This implies that $\lambda_{1}^{B}=g(g+n-1)$ is the first non-zero basic eigenvalue of $(S^{n},\mathcal{F})$.
The diameter of the leaf space $S^{n}/\mathcal{F}$ is exactly the distance between two singular leaves, which are level sets $V^{-1}(-1)$ and $V^{-1}(1)$ of the map $V=F|_{S^{n}}$, and hence we get $\mathrm{d}_{S^{n}/\mathcal{F}}=\pi/g$ (see Satz 4, [18]).
@@ 4. The case of Non-negative Ricci curvature
We recall again the gradient estimate by Zhong and Yang in Proposition 4.1, see [29] (also see the Theorem 2.3.2 in [6]). Once notices that their argument works with not the first eigenvalue, but also with an arbitrary eigenvalue $\lambda>0$ in the spectrum.
@@
Remark 4.1.
Notice that $\psi$ is an odd function which is continuous on $\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ and is the solution of the ODE
@@
See [24], section 44 of chapter 33, or also see [8].
We are now ready to prove Theorem 1.1. In the following proof, it is clear that when we have a gradient estimate, we can integrate over a geodesic joining two points of level sets of $u$ on both sides of the gradient estimate to derive an estimate for the first non-zero eigenvalue as many works in [11, 29] have done. The new thing in this proof is that we need to choose a more specific geodesic that works with the distances of leaves.
@@
First, we will recall a remarkable works of G.Xu and X.Xue that will be used in our main theorem of this section. More precisely, we will use Proposition 3.2 and a part that is modified from the statement of Theorem 4.1 (1), see [26]. Here we try to provide the outline of each proof, and for more detail, please visit [26], pages 8 - 13.
Proposition 4.2 (Proposition 3.2 (N) in [26]).
@@
In the next result of G.Xu and X.Xue that we want to use, the statement of Theorem 4.1 (1) in [26] does not seem to be correct as stated. More precisely, they claimed that with the same given conditions as in Proposition 4.2 and in addition, $\partial M=\emptyset$, then $M^{n}$ is isometric to $S^{1}\left(\frac{\mathcal{N}(u)}{2\sqrt{\lambda}}\right)\times N^{n-1}$, where $\mathcal{N}(u)$ is the number of nodal domains of $u$ and $S^{1}\left(\frac{\mathcal{N}(u)}{2\sqrt{\lambda}}\right)$ is a round circle with radius $\frac{\mathcal{N}(u)}{2\sqrt{\lambda}}$ and $\mathcal{N}(u)=2k$ for some $k\in\mathbb{Z}_{+}$. We provide a class of counterexamples for this statement: consider $M=\mathbb{R}\times_{\mathbb{Z}}N=\mathbb{R}\times N/\mathbb{Z}$ where $N^{n-1}$ is an (n1)(n-1)-dimensional compact Riemannian manifold of nonnegative Ricci curvature and the $\mathbb{Z}$-action is defined by $k\cdot(t,x)=\left(t+\frac{2k\pi}{\sqrt{\lambda}},\varphi^{k}(x)\right)$, where $k\in\mathbb{Z},(t,x)\in\mathbb{R}\times N$ and $\varphi:N\to N$ is an isometry on $N$. As mentioned earlier in Theorem 1.2, $M=\left[0,\frac{2\pi}{\sqrt{\lambda}}\right]\times N/\sim$ is the mapping torus of $\varphi$, where $(0,x)\sim\left(\frac{2\pi}{\sqrt{\lambda}},\varphi(x)\right)$. The function $u([(t,x)])=-\cos(\sqrt{\lambda}t)$ is the eigenfunction of $\lambda$ that satisfies the given conditions with the number of nodal domains $\mathcal{N}(u)=2$. In the case of a unit circle $N=S^{1}$, $\lambda=1$ and we obtain torus, Klein bottle (not a product) with the corresponding isometries $\varphi$ are the identity, the reflection respectively. We can even have many good examples of orientable manifolds that are not products by constructing mapping tori of some isometries, for example, we choose a torus $N=\mathbb{T}^{2}=S^{1}\times S^{1}=\mathbb{R}^{2}/\mathbb{Z}^{2}$ and construct a mapping torus of the isometry $\varphi=\left[\begin{matrix}-1&0\\ 0&-1\end{matrix}\right]$. This is one of the six compact, connected orientable flat Riemannian Seifert manifolds. Recall that a Seifert manifold is a closed 3-dimensional manifold together with a decomposition into a disjoint union of circles (called fibers) such that each fiber has a tubular neighborhood that forms a standard fibered torus. If the mapping torus $M_{\varphi}$ of $\varphi=\left[\begin{matrix}-1&0\\ 0&-1\end{matrix}\right]$ is a product of $S^{1}(r)\times P$ where $S^{1}(r)$ is a circle of some radius $r$ and $P$ is a compact $2$-dimensional manifold, then we imply that $P$ is flat. To make the product is orientable and connectedness, the only possibility of $P$ is to be a torus, and then, $S^{1}(r)\times P$ is also a compact, connected orientable flat Riemannian Seifert manifold which can be distinguished with $M_{\varphi}$ as classified. For more detail of the classification, see Theorem 2, section 8.2 in [21].
@@
Therefore, in the following theorem, we recall the given conditions from Theorem 4.1(1) of [26] with a modified conclusion and a modified argument for Step 3 of the proof of Theorem 4.1(1) in [26].
Theorem 4.3 (A modification of Theorem 4.1 (1) in [26]).
@@
Step 1 and Step 2 (see proof of Theorem 4.1 in [26]): Let $N^{n-1}=u^{-1}(-1)\cap\overline{M_{p}}$, which is an (n1)(n-1)-dim hypersurface in $M^{n}$ by Proposition 4.2 (if $u^{-1}(-1)\cap\overline{M_{p}}=\emptyset$, we define $N^{n-1}=u^{-1}(1)\cap\overline{M_{p}}$, similar argument applies); and $U_{\varepsilon}(N^{n-1})=\{x\in M^{n}:\mathrm{d}(x,N^{n-1})<\varepsilon\}$. Then we get a local chart $\{x_{1},x^{\prime}\}$ for $U_{r_{0}}(N^{n-1})$, where $x^{\prime}=(x_{2},\cdots,x_{n})$ is a coordinate chart for $N^{n-1}$ and
@@
Step 3: As claimed in [26], for any $q\in N^{n-1}$, we can find $z\in U_{r_{0}}^{-}(N^{n-1})$ such that $F(z)=F(q)=0$ and $F(z)=\max_{x\in\overline{U_{r_{0}}^{-}(N^{n-1})}}F(x)=0$. Note $|\nabla u(z)|>0,|u(z)|<1$ and $|\nabla\sin^{-1}u(z)|^{2}=\lambda$; from Proposition 4.2, we get the local splitting domain on $M_{z}$, a component of $\{x\in M:|\nabla u(x)|>0\}$ containing $z$.
@@
Remark 4.2.
In the case when $\mathcal{F}$ is given as the (connected) fibers of a Riemannian submersion $\pi:M\to M/\mathcal{F}$ with totally geodesic (or, more generally, minimal) fibers, if we apply Hang-Wang rigidity for the first non-zero eigenvalue of $M/\mathcal{F}$ which is also the first non-zero basic eigenvalue of $M$, then we can only conclude that $M/\mathcal{F}$ is isometric to $S^{1}\left(\frac{1}{\sqrt{\lambda_{1}^{B}}}\right)$ without knowing the structure of $\mathcal{F}$. By proving Theorem 1.2, not just $M/\mathcal{F}$, but we also understand the shape of $M$ and the structure of $\mathcal{F}$.